OpenStudy (anonymous):

A rock is dropped from the top of a 100 meter tall cliff? a)how long will it take for the rock to reach the ground? 4.51s b)What will be the velocity of the rock when it hits the ground? 44.27 m/s c) how far did the rock travel during (and only during) the third second of its motion? I figured part a and b. cant figure c out. please help and explain

OpenStudy (anonymous):

I understand that, but for some reason I get 44.1 m.

OpenStudy (anonymous):

sure lol

OpenStudy (anonymous):

\[v _{2}^2 = 2(-9.8m/s^2)(y2)\]

OpenStudy (anonymous):

the equation i used. I assumed it started at 0m

OpenStudy (anonymous):

haha by common sense its 58.8m, but I dont know how to get it in formula form...

OpenStudy (shane_b):

Your first two answers are correct. For the third it's just asking how far it traveled between the 2s and 3s mark...which will be the difference in distance at those two times: \[\large d=V_0t+\frac{1}{2}at^2\]\[\large \Delta d=\frac{1}{2}(-9.8m/s^2)(3s)^2-\frac{1}{2}(-9.8m/s^2)(2s)^2=-24.5m\]